Delph Company uses a job-order costing system and has two manufacturing departments-Molding and Fabrication. The company provided the following estimates at the beginning of the year:


MoldingFabricationTotal
Machine-hours26,00036,00062,000
Fixed manufacturing overhead costs$ 780,000$200,000$980,000
Variable manufacturing overhead cost per machine-hour$ 5.40$ 5.40

During the year, the company had no beginning or ending inventories and it started, completed, and sold only two jobs-Job D-70 and Job C-200. It provided the following information related to those two jobs:

Job D-70:MoldingFabricationTotal
Direct materials cost$374,000$322,000$696,000
Direct labor cost$220,000$170,000$390,000
Machine-hours17,0009,00026,000

Job C-200:MoldingFabricationTotal
Direct materials cost $210,000$210,000$210,000$420,000
Direct labor cost$170,000$290,000$460,000
Machine-hours9,00027,00036,000

Delph had no underapplied or overapplied manufacturing overhead during the year.

2. Assume Delph uses departmental predetermined overhead rates based on machine-hours.

a. Compute the departmental predetermined overhead rates.

b. Compute the total manufacturing cost assigned to Job D-70 and Job C-200.

C. If Delph establishes bid prices that are 140 % of total manufacturing costs, what bid prices would it have established for Job D-70 and Job C-200?

d. What is Delph's cost of goods sold for the year?
Answer #1

2) Departmental overhead rate :

Molding overhead rate = (780000/26000)+5.40 = 35.40 per Hour

Fabrication overhead rate = (200000/36000)+5.40 = 10.96 per hour

2b) Total manufacturing cost

Job D-70 Job C-200
Direct material 696000 420000
Direct labor 390000 460000
Molding overhead 601800 318600
Fabrication overhead 98640 295920
Total manufacturing cost 1786440 1494520

2c) Bid price

Job D-70 = 1786440*1.4 = 2501016

Job C-200 = 1494520*1.4 = 2092328

2d) Cost of goods sold = 1786440+1494520 = 3280960